OpenStudy (marigirl):

Could someone please help me understand this question and specifically what it means by proportional: Water leaks from the bottom of a tank at a rate proportional to the depth h of the water in the tank. Write down an equation describing this if V is the volume of the water in the tank after time 't' and if the tank is a cylinder show that dh/dt is proportional to h.

OpenStudy (marigirl):

@IrishBoy123

OpenStudy (marigirl):

@AlexandervonHumboldt2

OpenStudy (marigirl):

so far i understand that \[\frac{ dV }{ dt }\] which is the water leaking is proportional to the depth...How can i mathematically express this?

OpenStudy (irishboy123):

\[\frac{ dV }{ dt } \propto h\] means \[\frac{ dV }{ dt } = k\; h\] where k is some constant now start processing this with \(V = \pi r^2 h\) in mind

OpenStudy (marigirl):

\[\frac{ dV }{ dt }=\pi r^2 \frac{ dh }{ dt }\] so then right side will equal kh

OpenStudy (irishboy123):

no, LHS will be kh because \(\frac{ dV }{ dt } = k\; h\)

OpenStudy (marigirl):

\[so \frac{ dh }{ dt }=\frac{ kh }{ \pi r^2 }\]

OpenStudy (irishboy123):

yes \[so \frac{ dh }{ dt }=\frac{ kh }{ \pi r^2 } = const \times h\]

OpenStudy (irishboy123):

or \[\frac{dh}{dt} \propto h\]

OpenStudy (marigirl):

ah yes! which is what we discussed at the beginning!! thanks! in a nutshell - can you explain what it means when they say proportional

OpenStudy (irishboy123):

for proportional think straight lines, eg \(y = mx + x\), which means \(y \propto x\) the relationship is linear, as x increases y increases. contrast with inversely proportional. here think \(y = \frac{1}{x}\), so \(y \propto \frac{1}{x}\), as y is "inversely proportional" to x

OpenStudy (marigirl):

Great ! thanks heaps!